OpenStudy (anonymous):

I'm doing the intro to programming/python, if anyone else is running through it and would like to share some code/examples/problem, my email is snibor.ds@gmail.com Thanks

OpenStudy (anonymous):

I think this works: ps1.py def isPrime(n): # this is a function created by me ri = 2 rf = n / 2 isPrim = 1 # yes for b in range(ri, rf+1): mod = n % b if mod == 0: isPrim = 0 # no return isPrim x = 2 lastPrime = 0 numbPrime = 1000 while lastPrime < numbPrime: # range from 0 to 999, i.e. 1000 if isPrime(x): # call function isPrime() print x lastPrime = lastPrime + 1 x = x + 1 Try: c:\Python27\ps1.py > ps1.txt to see the results in a .txt file.